# Lyapunov Exponents versus Integrability in Random Conservative Dynamics

### v1

@@ Abstract.
We consider random dynamical systems generated by volume-preserving piecewise $C^{1}$ maps. For this class of random systems, we establish an invariance principle asserting that if the Lyapunov exponents vanish, then there exists a measurable family of probability measures on the projective bundle that is invariant under the projective cocycle induced by the derivative. We apply this principle to two classes of random systems. First, we study random additive perturbations of a single billiard map associated with a strictly convex planar table on a surface of constant curvature. In this setting, we prove that the Lyapunov exponents vanish almost everywhere if and only if the billiard table is a geodesic disk. Second, we consider random additive perturbations of a single standard map and show that the Lyapunov exponents vanish almost everywhere if and only if the standard map is integrable.
@@ 1. Introduction
Numerical simulations indicate that generic conservative surface maps exhibit a mixed phase space, with coexisting regions of regular and chaotic dynamics. This behavior is observed in a broad class of systems, including area-preserving twist maps and billiards in convex tables, and suggests the presence of positive metric entropy with respect to the Lebesgue measure. Despite substantial numerical and heuristic evidence, no general method is currently available to establish this property rigorously.
In this paper, we consider random additive perturbations of area-preserving maps on the 2-torus ${\mathbb{T}}^{2}$ and investigate whether randomness leads to non-vanishing Lyapunov exponents almost everywhere. More precisely, given an area-preserving map $f:\mathbb{T}^{2}\to\mathbb{T}^{2}$, for each $x\in\mathbb{T}^{2}$, we define the corresponding additive perturbation by
@@
We consider random compositions generated by an i.i.d. sequence $(X_{n})_{n\geq 0}$ of $\mathbb{T}^{2}$-valued random variables whose support contains a neighborhood of 0. The resulting random dynamical system is given by
@@
Since each map $f_{x}$ preserves Lebesgue measure, this construction defines an area-preserving random dynamical system on $\mathbb{T}^{2}$.
The introduction of additive randomness is expected to weaken invariant structures and enhance ergodic properties, increasing the likelihood of exponential instability and positive metric entropy.
This leads to the following question: do conservative random perturbations typically generate non-vanishing Lyapunov exponents almost everywhere, and if not, does the persistence of vanishing Lyapunov exponents characterize integrability of the underlying deterministic system?
@@
The first class consists of convex billiards on surfaces of constant curvature. Let $S$ be a Riemannian surface of constant curvature, and let $D\subset S$ be a strictly convex domain whose boundary $\partial D$ is a $C^{2}$ simple closed curve with positive geodesic curvature. We refer to such a domain $D$ as a convex table. A billiard in $D$ describes the motion of a point particle traveling along geodesics of $S$ inside $D$ and undergoing elastic reflections at $\partial D$.
The dynamics is encoded by the billiard map $\Phi$, an area-preserving twist map defined on the space of collisions with the boundary. Using arc-length coordinates along $\partial D$ and the reflection angle, the collision space can be identified with $V=S^{1}\times[-1,1]$. It is well known that $\Phi$ is an area-preserving twist map preserving the Lebesgue measure on $V$. Identifying the collision space with the $2$-torus ${\mathbb{T}}^{2}$, we denote by $f\colon{\mathbb{T}}^{2}\to{\mathbb{T}}^{2}$ the map induced by $\Phi$. Full details of this construction are given in Section 3.
Despite extensive study, it remains an open problem whether billiards in strictly convex tables with smooth boundary can exhibit positive metric entropy. All known examples of billiards with positive metric entropy involve tables that are either not strictly convex or not smooth; see, for instance [11, 33, 15, 27, 22]. Although the topological entropy of convex billiards is generically positive, this does not imply positivity of metric entropy.
Additive perturbations $f_{x}$ of the billiard map $f$ have a transparent geometric interpretation. Given $x=(\bar{s},\bar{r})$ and a collision $(s_{0},r_{0})$, one computes the next collision $(s_{1},r_{1})=f(s_{0},r_{0})$, and then applies a translation, shifting the arc-length coordinate by $\bar{s}$ modulo $L$ and the $r$-coordinate by $\bar{r}$ modulo $2$.
Our first main result characterizes precisely when Lyapunov exponents vanish for random billiards.
@@
Let $D$ be a convex billiard table on a surface $S$ of constant curvature, and let $(X_{n})_{n\geq 0}$ be an i.i.d. sequence of random variables taking values in $\mathbb{T}^{2}$ with distribution that is absolutely continuous with respect to the Lebesgue measure on $\mathbb{T}^{2}$ and whose support contains a nonempty open set. Then the Lyapunov exponents of the associated random dynamical system $\{\mathsf{T}^{n}\}$ vanish almost everywhere if and only if $D$ is a geodesic disk.
This theorem reflects the exceptional symmetry of circular billiards on surfaces of constant curvature. For such billiards, the phase space of the billiard map $f$ is foliated by non-contractible invariant circles [7, 16]. This property is sometimes referred to as total integrability. When $f$ is replaced by a conservative random perturbation $f_{x}$, obtained by composing $f$ with a translation, this foliation remains globally invariant: although individual circles are no longer invariant under each realization $f_{x}$, the map $f_{x}$ sends each circle of the foliation onto another circle. As a consequence, the random dynamics exhibits no exponential growth of derivatives.
This result significantly extends the work of Nguyen, who showed that for billiards with non-circular elliptical tables in the Euclidean plane, conservative random perturbations have non-vanishing Lyapunov exponents almost everywhere [28].
The second class of conservative dynamical systems considered in this paper consists of standard maps and their generalizations. The standard map is a family of diffeomorphisms $f\colon\mathbb{T}^{2}\to\mathbb{T}^{2}$ depending on a real parameter $K\in\mathbb{R}$, defined by
@@
Standard maps arise naturally in Hamiltonian dynamics as models of periodically kicked systems and play a paradigmatic role in the study of the transition from integrability to chaos. However, proving the existence of non-vanishing Lyapunov exponents for non-integrable standard maps and their generalizations remains notoriously difficult.
Our second main result establishes a rigidity theorem for random additive perturbations of standard maps analogous to Theorem 1.1. In this case we are able to get the result for a larger class of measures. We say that a probability measure $\nu$ on $\mathbb{T}^{2}$ is non-degenerate if it satisfies one of the following conditions:
@@
    $\nu$ is absolutely continuous with respect to the Lebesgue measure on $\mathbb{T}^{2}$ and its support contains a nonempty open set;
@@
Let $f\colon\mathbb{T}^{2}\to\mathbb{T}^{2}$ be a standard map, and let $(X_{n})_{n\geq 0}$ be an i.i.d. sequence of $\mathbb{T}^{2}$-valued random variables with non-degenerate distribution. Then the Lyapunov exponents of the associated random dynamical system $\{\mathsf{T}^{n}\}$ vanish almost everywhere if and only if $f$ is integrable, that is, if and only if $K=0$.
For random standard maps, Blumenthal, Xue and Young obtained quantitative lower bounds on the maximal Lyapunov exponent under sufficiently strong random additive perturbations; see [9]. By contrast, the present work does not rely on large randomness and instead provides a sharp characterization of when Lyapunov exponents vanish.
Taken together, Theorems 1.1 and 1.2 show that for conservative random perturbations of billiards and standard maps, vanishing Lyapunov exponents occur if and only if the underlying deterministic system is totally integrable.
The conceptual core of the paper is a rigidity result for random dynamical systems, which we call an invariance principle following the terminology of Avila–Viana [2]. Roughly speaking, we prove that if the Lyapunov exponents of a random system coincide almost everywhere, then there exists a non-random invariant structure on the projective tangent bundle. This result is stated precisely in Theorem 2.10 and is obtained by applying a theorem of Ledrappier [26] to the setting of random maps. This rigidity result is rooted in Furstenberg’s theory of random matrix products [20] and its subsequent extensions by Ledrappier, Baxendale, and Avila–Viana.
The paper is organized as follows. Section 2 establishes the general invariance principle for random maps. Section 3 provides background on convex billiards on surfaces of constant curvature. In Section 3.3, we prove Theorem 1.1. Finally, Section 4 is devoted to random standard maps and contains the proof of Theorem 1.2.
@@ 2.3.1. Ledrappier’s Theorem
Let $(\Omega^{\prime},\Sigma^{\prime})$ be the two-sided product space $\Omega^{\prime}=Z^{{\mathbb{Z}}}$ endowed with the product $\sigma$-algebra $\Sigma^{\prime}:={\mathcal{Z}}^{{\mathbb{Z}}}$. For each $n\in{\mathbb{Z}}$, let $\pi_{n}\colon\Omega^{\prime}\to Z$ be the projection given by $\pi_{n}(\omega^{\prime})=\omega^{\prime}_{n}$ for all $\omega^{\prime}=(\omega^{\prime}_{n})_{n\in{\mathbb{Z}}}\in\Omega^{\prime}$. Since $Z$ is a separable complete metric space, there exists a probability measure ${\mathbb{P}}_{\zeta}$ on $\Sigma^{\prime}$ such that the coordinate process $\{\pi_{n}\}_{n\in{\mathbb{Z}}}$ is a Markov chain on $(\Omega^{\prime},\Sigma^{\prime},{\mathbb{P}}_{\zeta})$ with transition probability $P_{z}$ and stationary probability measure $\zeta$. Moreover, the shift map $\sigma^{\prime}\colon\Omega^{\prime}\to\Omega^{\prime}$ is measurable with respect to $\Sigma^{\prime}$ and preserves ${\mathbb{P}}_{\zeta}$ (see [18, Theorem 3.1.7 and Example 5.1.3]).
@@ 2.3.2. Invariant principle for random maps
The argument below follows closely the proof of [12, Theorem 1].
@@ 2.3.3. When $\nu$ is absolutely continuous
The final part of the theorem follows from (2.2) and standard results on linear cocycles [30, Sections 6.3 and 7.3]. ∎
@@
When $d=2$, the last part of Proposition 2.12 reduces to the following characterization (see [30, Section 6.3]):
@@
The failure of Conditions (1) and (2) in the conclusion of Proposition 2.12 implies that $\lambda^{+}_{F}>0$ a.e. on $\Omega\times{\mathbb{T}}^{d}$. This provides a criterion for the positivity of the maximal Lyapunov exponent of $F$ analogous to Furstenberg’s criterion for random matrices [30, Section 7.3].
@@ 2.4. Some ergodic properties of random perturbations of continuous toral maps
Since $\{P_{y}\}$ is strong Feller by Lemma 2.18 and the support of $m$ is the whole ${\mathbb{T}}^{d}$, it follows from [4, Part (iii) of Proposition 5.18] that $\{P_{y}\}$ is uniquely ergodic. Consequently, $m$ is an ergodic stationary probability measure for $\{P_{y}\}$. Finally, by [30, Proposition 5.13], the invariant probability measure $\rho_{\nu}\times m$ of $F$ is ergodic. ∎
@@
By Breiman’s Law of Large Numbers for Markov chains [5, Corollary 2.7], applied to the uniquely ergodic kernel $P_{y}$, the equidistribution property holds; see also [20, Lemma 7.1].
@@ 3.1.1. Billiard map in coordinates $(s,\theta)$
Moreover, in all cases, the derivative $D\phi$ admits the following limits [24] and [14, Proposition 3.4]: for every $s\in S^{1}$,

### v2

# Positive Lyapunov Exponents versus Integrability in Random Conservative Dynamics

@@ Abstract.
We study random dynamical systems generated by volume-preserving piecewise $C^{1}$ maps. For this class of systems, we establish an invariance principle stating that if all Lyapunov exponents vanish, then there exists a measurable family of probability measures on the projective bundle that is invariant under the projective cocycle induced by the derivative. We apply this principle to two classes of random systems. First, we consider random additive perturbations of the billiard map associated with a strictly convex planar table on a surface of constant curvature. In this setting, we show that the Lyapunov exponents vanish almost everywhere if and only if the billiard table is a geodesic disk. Second, we study random additive perturbations of a standard map and prove that the Lyapunov exponents vanish almost everywhere if and only if the map is integrable.
@@ 1. Introduction
Numerical simulations indicate that generic area-preserving surface maps exhibit a mixed phase space, with coexisting regions of regular and chaotic dynamics. This behavior is observed in a broad class of systems, including area-preserving twist maps and billiards in convex tables, and suggests positivity of the metric entropy with respect to the Lebesgue measure. Despite substantial numerical and heuristic evidence, no general method is currently available to establish this property rigorously.
In this paper, we study random additive perturbations of area-preserving maps on the 2-torus ${\mathbb{T}}^{2}$ and investigate whether randomness leads to non-vanishing Lyapunov exponents almost everywhere. More precisely, given an area-preserving map $f:\mathbb{T}^{2}\to\mathbb{T}^{2}$, for each $x\in\mathbb{T}^{2}$, we define the corresponding additive perturbation by
@@
We then consider random compositions generated by an i.i.d. sequence $(X_{n})_{n\geq 0}$ of $\mathbb{T}^{2}$-valued random variables whose support contains a neighborhood of the origin. The resulting random dynamical system is given by
@@
Since each map $f_{x}$ is area-preserving, this construction yields an area-preserving random dynamical system on $\mathbb{T}^{2}$.
The introduction of additive randomness is expected to weaken invariant structures and enhance ergodic properties, increasing the likelihood of exponential instability and positive metric entropy. This leads to the following question: do conservative random perturbations typically generate non-vanishing Lyapunov exponents almost everywhere, and if not, does the persistence of vanishing Lyapunov exponents characterize integrability of the underlying deterministic system?
@@
We begin with billiards on surfaces of constant curvature. Let $S$ be a Riemannian surface of constant curvature, and let $D\subset S$ be a strictly convex domain whose boundary $\partial D$ is a $C^{2}$ simple closed curve with positive geodesic curvature. We refer to such a domain $D$ as a convex table. A billiard in $D$ describes the motion of a point particle traveling along geodesics of $S$ inside $D$ and undergoing elastic reflections at $\partial D$.
The dynamics is encoded by the billiard map $\Phi:V\to V$, a twist map defined on the collision space $V=S^{1}\times[-1,1]$. Let $L$ denote the length of $\partial D$. Each point $(s,r)\in V$ uniquely represents a collision with the boundary: the coordinate $s\in\mathbb{R}/(L\mathbb{Z})\cong S^{1}$ is the arc-length parameter of the collision point on $\partial D$, while $r=-\cos\theta\in[-1,1]$, where $\theta\in[0,\pi]$ is the reflection angle measured with respect to the tangent to $\partial D$ at $s$. The map $\Phi$ sends each collision $(s,r)$ to the next collision along the trajectory. Upon identifying $V$ with the two-torus ${\mathbb{T}}^{2}$, the billiard map $\Phi$ induces a map $f:{\mathbb{T}}^{2}\to{\mathbb{T}}^{2}$ that preserves the Lebesgue measure. Full details of this construction are given in Section 3.
Despite extensive study, it remains an open problem whether billiards in strictly convex tables with smooth boundary can exhibit positive metric entropy with respect to the Lebesgue measure. All known examples of billiards with this property involve tables that are either not strictly convex or not smooth; see, for instance, [12, 36, 18, 30, 25]. It is worth mentioning that, while generic convex billiards have positive topological entropy [7, 14, 37], this fact alone does not yield examples with positive metric entropy with respect to the Lebesgue measure.
Our first main result characterizes precisely when Lyapunov exponents vanish for random billiards, where $f$ denotes the billiard map on ${\mathbb{T}}^{2}$, $f_{x}$ its additive perturbations, and $\{\mathsf{T}^{n}\}$ the associated random dynamical system. Additive perturbations $f_{x}$ of the billiard map $f$ admit a clear geometric interpretation. Given $x=(\bar{s},\bar{r})$ and an initial collision $(s_{0},r_{0})$, one first computes the next collision $(s_{1},r_{1})=f(s_{0},r_{0})$, and then translates the collision point by $x$ in the phase space, shifting the arc-length coordinate by $\bar{s}$ modulo $L$ and the $r$-coordinate by $\bar{r}$ modulo $2$.
@@
Let $D$ be a convex billiard table on a surface $S$ of constant curvature. Let $(X_{n})_{n\geq 0}$ be an i.i.d. sequence of $\mathbb{T}^{2}$-valued random variables with a distribution that is absolutely continuous with respect to the Lebesgue measure on ${\mathbb{T}}^{2}$ and has support containing a neighbourhood of the origin. Then the Lyapunov exponents of the associated random dynamical system $\{\mathsf{T}^{n}\}$ vanish almost everywhere if and only if $D$ is a geodesic disk.
This theorem reflects the exceptional symmetry of circular billiards on surfaces of constant curvature. For such billiards, the phase space of the map $f$ is foliated by non-contractible invariant circles [8, 19], a property often referred to as total integrability. When $f$ is replaced by an additive perturbation $f_{x}$, this foliation remains globally invariant: although individual circles may no longer be invariant under $f_{x}$, each circle of the foliation is mapped by $f_{x}$ onto another circle. As a consequence, the associated random dynamical system exhibits no exponential growth of derivatives, and hence has vanishing Lyapunov exponents almost everywhere.
Theorem 1.1 extends significantly the work of Nguyen, who showed that for billiards in non-circular elliptical tables in the Euclidean plane, random additive perturbations have non-vanishing Lyapunov exponents almost everywhere [31].
The second class of conservative dynamical systems considered in this paper consists of standard maps and their generalizations. The standard map, introduced by Chirikov [16], is a family of diffeomorphisms $f\colon\mathbb{T}^{2}\to\mathbb{T}^{2}$ depending on a real parameter $K\in\mathbb{R}$, defined by
@@
Standard maps arise naturally in Hamiltonian dynamics as models of periodically kicked systems and play a central role in the study of the transition from integrability to chaos. However, proving the existence of non-vanishing Lyapunov exponents for non-integrable standard maps and their generalizations remains notoriously difficult.
Our second main result establishes a rigidity theorem for random additive perturbations of standard maps, analogous to Theorem 1.1. In this setting, the conclusion holds under weaker assumptions on the common distribution of the random variables $(X_{n})_{n\geq 0}$. We say that a probability measure $\nu$ on $\mathbb{T}^{2}$ is non-degenerate if it satisfies one of the following conditions:
@@
    $\nu$ is absolutely continuous with respect to the Lebesgue measure on $\mathbb{T}^{2}$ and its support contains a neighbourhood of the origin;
@@
Let $f\colon\mathbb{T}^{2}\to\mathbb{T}^{2}$ be a standard map, and let $(X_{n})_{n\geq 0}$ be an i.i.d. sequence of ${\mathbb{T}}^{2}$-valued random variables with a non-degenerate distribution. Then the Lyapunov exponents of the associated random dynamical system $\{\mathsf{T}^{n}\}$ vanish almost everywhere if and only if $f$ is integrable, that is, if and only if $K=0$.
For random standard maps, Blumenthal, Xue, and Young obtained quantitative lower bounds on the maximal Lyapunov exponent under sufficiently strong random additive perturbations [10]. In contrast, the present work does not rely on large randomness and instead provides a sharp characterization of the vanishing of Lyapunov exponents.
Taken together, Theorems 1.1 and 1.2 show that for random conservative perturbations of billiards and standard maps, vanishing Lyapunov exponents occur if and only if the underlying deterministic system is totally integrable.
The conceptual core of the paper is a rigidity result for random dynamical systems, which we call an invariance principle, following the terminology of Avila–Viana [2]. Roughly speaking, we show that if the Lyapunov exponents of a random system coincide almost everywhere, then there exists a non-random invariant structure on the projective tangent bundle. This is stated precisely in Theorem 2.10 and is derived by applying a theorem of Ledrappier [29] to the setting of random maps considered in this paper. The argument is rooted in Furstenberg’s theory of random matrix products [23] and its subsequent extensions to stationary sequences of matrices by Ledrappier [29] and to random diffeomorphisms on manifolds by Carverhill [13] and Baxendale [3].
The paper is organized as follows. Section 2 establishes the general invariance principle for random maps and concludes with results on ergodic properties of the random dynamical systems considered here. Section 3 provides background on convex billiards on surfaces of constant curvature. The proof of Theorem 1.1 is given in Section 3.3. Finally, Section 4 is devoted to random standard maps and contains the proof of Theorem 1.2.
@@ 2.3.1. Ledrappier’s Theorem
Let $(\Omega^{\prime},\Sigma^{\prime})$ be the two-sided product space $\Omega^{\prime}=Z^{{\mathbb{Z}}}$ endowed with the product $\sigma$-algebra $\Sigma^{\prime}:={\mathcal{Z}}^{{\mathbb{Z}}}$. For each $n\in{\mathbb{Z}}$, let $\pi_{n}\colon\Omega^{\prime}\to Z$ be the projection given by $\pi_{n}(\omega^{\prime})=\omega^{\prime}_{n}$ for all $\omega^{\prime}=(\omega^{\prime}_{n})_{n\in{\mathbb{Z}}}\in\Omega^{\prime}$. Since $Z$ is a separable complete metric space, there exists a probability measure ${\mathbb{P}}_{\zeta}$ on $\Sigma^{\prime}$ such that the coordinate process $\{\pi_{n}\}_{n\in{\mathbb{Z}}}$ is a Markov chain on $(\Omega^{\prime},\Sigma^{\prime},{\mathbb{P}}_{\zeta})$ with transition probability $P_{z}$ and stationary probability measure $\zeta$. Moreover, the shift map $\sigma^{\prime}\colon\Omega^{\prime}\to\Omega^{\prime}$ is measurable with respect to $\Sigma^{\prime}$ and preserves ${\mathbb{P}}_{\zeta}$ (see [21, Theorem 3.1.7 and Example 5.1.3]).
@@ 2.3.2. Invariant principle for random maps
The argument below follows closely the proof of [13, Theorem 1].
@@ 2.3.3. When $\nu$ is absolutely continuous
The final part of the theorem follows from (2.2) and standard results on linear cocycles [33, Sections 6.3 and 7.3]. ∎
@@
When $d=2$, the last part of Proposition 2.12 reduces to the following characterization (see [33, Section 6.3]):
@@
The failure of Conditions (1) and (2) in the conclusion of Proposition 2.12 implies that $\lambda^{+}_{F}>0$ a.e. on $\Omega\times{\mathbb{T}}^{d}$. This provides a criterion for the positivity of the maximal Lyapunov exponent of $F$ analogous to Furstenberg’s criterion for random matrices [33, Section 7.3].
@@ 2.4. Some ergodic properties of random perturbations of continuous toral maps
Since $\{P_{y}\}$ is strong Feller by Lemma 2.18 and the support of $m$ is the whole ${\mathbb{T}}^{d}$, it follows from [4, Part (iii) of Proposition 5.18] that $\{P_{y}\}$ is uniquely ergodic. Consequently, $m$ is an ergodic stationary probability measure for $\{P_{y}\}$. Finally, by [33, Proposition 5.13], the invariant probability measure $\rho_{\nu}\times m$ of $F$ is ergodic. ∎
@@
By Breiman’s Law of Large Numbers for Markov chains [5, Corollary 2.7], applied to the uniquely ergodic kernel $P_{y}$, the equidistribution property holds; see also [23, Lemma 7.1].
@@ 3.1.1. Billiard map in coordinates $(s,\theta)$
Moreover, in all cases, the derivative $D\phi$ admits the following limits [27] and [17, Proposition 3.4]: for every $s\in S^{1}$,